OpenStudy (anonymous):

Easy Medal In order to maintain a healthy weight, an average adult dog that weighs 16 kilograms must consume 12 grams of dog food per kilogram of its body weight each day. A specialty pet store sells dog food by the cup, where 6 cups of dog food weighs 606 grams. About how many cups of dog food should the dog consume each day in order to maintain a healthy weight? 2.00 0.53 0.32 1.90

OpenStudy (anonymous):

i figured this would be C because the amount of it being 6 is way over how many grams they need

jigglypuff314 (jigglypuff314):

I got D 12 * 16 = 192 606/6 = 101 192/101 = 1.90

OpenStudy (anonymous):

oh so basically multiplying the diving making it round off to 1.90

OpenStudy (anonymous):

thanks

jigglypuff314 (jigglypuff314):

you're welcome :)